OpenStudy (anonymous):

Quadratic formula. I can't seem to figure it out. I'll become a fan and give a medal

zepdrix (zepdrix):

Do you have a specific example that we need to use the Formula on? :)

OpenStudy (anonymous):

hold on i'm attaching a file

zepdrix (zepdrix):

\[\large\rm \color{green}{1}x^2+\color{orangered}{9}x+\color{royalblue}{10}=0\]And our formula is,\[\large\rm x=\frac{-\color{orangered}{b}\pm\sqrt{\color{orangered}{b}^2-4\color{green}{a}\color{royalblue}{c}}}{2\color{green}{a}}\]ya? :o Understand why I put a 1 on the x^2?

OpenStudy (anonymous):

nope. I suck at geometry

OpenStudy (anonymous):

A = 1 B = 9 C = 10

OpenStudy (anonymous):

also remember that if B is a negative i.e -9 then it would be +b rather than -b at the start

OpenStudy (anonymous):

\[\frac{-9\pm \sqrt{9^2-4*1*10} }{2*1 }\]

OpenStudy (anonymous):

\[X _{1} = \frac{ -9+\sqrt{41} }{2 } X _{2} = \frac{ -9-\sqrt{41} }{ 2 }\]

OpenStudy (anonymous):

Tbh @Lily2913 I don't understand any of that

OpenStudy (anonymous):

ok it's B

OpenStudy (anonymous):

-7.7, -1.3

OpenStudy (anonymous):

Thank you so so so so much. I've literally been working on that problem for an hour

OpenStudy (anonymous):

Quadratic formula gives you 2 results X1 and X2 so if you had put the above in your calculator you would have seen -7.7 and -1.3

OpenStudy (anonymous):

I hope you get a hang of it before any exams you need to take that use it. :)

OpenStudy (anonymous):

IKR currently i'm screwed